# Dynamic Financial Analysis (DFA) of General Insurers under Climate Change

### v1

@@ 1.1 Background
Climate change poses a multifaceted and profound impact on the financial performance and position of general insurers. On the liabilities side, shifting weather patterns are likely to alter the frequency and severity of future claims (see, e.g., Haug et al., 2011; Lyubchich et al., 2019), which could lead to increased claims costs and adversely affect underwriting profitability. On the assets side, climate change-induced fluctuations in temperature, shifting patterns of natural disasters, and evolving socio-economic conditions may significantly influence key macroeconomic variables. These include inflation rates (see, e.g., Parker, 2018; Xepapadeas and Economides, 2018; Kotz et al., 2024), interest rates (see, e.g., Bylund and Jonsson, 2020; Mongelli et al., 2022), and equity returns (see, e.g., Karydas and Xepapadeas, 2022; Venturini, 2022; Barnett, 2023), thereby impacting the investment performance of general insurers. Ultimately, the combined effects of climate change on both assets and liabilities will cascade into insurers’ capital positions, potentially affecting the financial health and stability of the broader general insurance market.
@@
To navigate these challenges, a holistic understanding of climate change’s impact on general insurers is essential. This need is increasingly recognised by regulators, as evidenced by the growing number of climate-related disclosure requirements, such as IFRS S2 (IFRS, 2023) and its Australian equivalent, AASB S2 (AASB, 2023). In the existing literature, several studies have investigated the impacts of climate change on individual financial items. For example, on the liabilities side, various studies have examined climate change effects on hazards such as floods (see, e.g., Seneviratne et al., 2021), bushfires (see, e.g., Quilcaille et al., 2022), and tropical cyclones and storms (see, e.g., Jagger et al., 2008, 2011; Meiler et al., 2022), generally indicating that rising temperatures exacerbate the frequency and severity of most hazard events (Chapter 11 of IPCC, 2021a, pp. 1546–1600). Additionally, the impact of climate change on non-catastrophe, weather-related claims frequency and severity have also been investigated in literature (see, e.g., Haug et al., 2011; Lyubchich et al., 2017). On the assets side, climate change impacts on interest rates, inflation rates, and equity returns have been explored both theoretically using General Equilibrium Models (Xepapadeas and Economides, 2018; Karydas and Xepapadeas, 2022; Barnett, 2023) and empirically using multiple regression (see, e.g., Parker, 2018) and Factor Models (see, e.g., Venturini, 2022). While efforts have been made to incorporate climate change impacts into the modelling of insurance liabilities and assets in various special (and limited) cases, as yet a holistic approach to assess the interconnected financial impacts of climate change on general insurers across both assets and liabilities remains lacking.
@@ 1.2 Statement of contributions
    Long-term nature of climate change: Climate change is inherently long-term, and it is usually challenging to distinguish the effects of global warming from internal variability over short horizons; indeed, internal variability ranges from inter-annual (e.g., El Niño–Southern Oscillation) to inter-decadal (e.g., Atlantic Multi-decadal Variability) (Chapter 11 of IPCC, 2021a, pp. 1546–1600). In addition, new regulations – such as IFRS S2 (IFRS, 2023) and AASB S2 (AASB, 2023) – require insurers to disclose the long-term financial impacts of climate change. A longer-term perspective is also essential for decisions by insurers and policymakers that extend across multiple years, including relocation planning (typically 8 to 40 years Bower and Weerasinghe, 2021) and reinsurance planning (influenced by cycles of about 9 to 11 years Meier and Outreville, 2006, 2010).
    An initial attempt to measure climate change effects on both assets and liabilities of general insurers simultaneously is presented in Gatzert and Özdil (2024), albeit limited to a one-year horizon. Our proposed framework utilizes a multi-year perspective within a DFA structure to assess the cumulative effects of climate change on general insurers, effectively capturing its long-term nature.
@@
    To address these gaps, we embed stochastic simulations within climate scenario analysis. Our framework captures uncertainty from both societal developments and the inherent randomness in natural and financial systems. By enabling simulation of a distribution of financial outcomes across diverse climate pathways, the framework supports assessments of both expected returns and risks, which are key considerations for insurers, actuaries, and regulators. While our analysis incorporates the economic and environmental assumptions underlying each scenario, evaluating the validity of those assumptions lies beyond the scope of this paper.
Our proposed framework is tailored to the general insurance sector by incorporating its unique features: it captures asset–liability interdependence through an interconnected structure and models the high variability of liability cash flows via DFA’s stochastic simulations. It also includes catastrophe reinsurance programs and accounts for the sensitivity of reinsurance premiums to capital constraints, which are also key factors affecting insurer profitability and solvency (see, e.g., Meier and Outreville, 2006).
@@ 2.1 Model overview
The proposed climate-dependent DFA framework is illustrated in Figure 2.1. A key input to this framework is the set of climate scenarios, for which we adopt the Shared Socioeconomic Pathways (SSPs). These SSPs form a widely adopted framework in climate research and they are central to the IPCC’s climate risk assessments (O Neill et al., 2017). Each SSP scenario is associated with a narrative, from which the economic growth rate at the technological frontier is derived (Dellink et al., 2017a). Starting from historical values, country-specific GDP projections are generated under the assumption that individual economies gradually converge toward this frontier. The convergence speed is determined by the degree of trade openness, as inferred from the scenario narratives (Dellink et al., 2017a). The emissions pathways consistent with the economic and environmental assumptions underlying each scenario are then used as inputs to climate models to produce projections of future climate at a much finer spatial resolution, typically at the level of gridded cells (Eyring et al., 2016).
@@
    SSP 4.5 (“Middle of the Road”): Represents a development pathway aligned with typical historical trends observed over the past century, leading to moderate mitigation and adaptation challenges.
@@
Note that caution is warranted when interpreting the results derived from these scenarios, given the limitations of their underlying assumptions, particularly under high-emission pathways such as SSP 8.5. This scenario assumes continued economic growth without accounting for the risk of economic collapse (Keen et al., 2021). Moreover, labour productivity is treated as exogenous in SSP scenarios, overlooking potential negative impacts of climate change (Keen et al., 2021). While SSP 8.5 also assumes strong investments in health, education, and highly engineered infrastructure (O Neill et al., 2017), which may mitigate climate-related productivity losses, the extent of such mitigation remains uncertain. Assessing or refining these assumptions lies beyond the scope of this paper and is left for future research.
@@
This paper focuses on the direct financial impacts of climate change on general insurers, as outlined in the modelling framework presented in Figure 2.1. While indirect effects, such as shifts in customer preference toward “greener” insurers or the potential rise in liability risks (e.g., lawsuits against commercial policyholders for environmental damage Alien, 2003; Bullock, 2022) are not included here, these are important areas for future research. In addition, we do not consider unidentified risks stemming from unforeseen responses of environmental and social systems to climate change, as these currently remain entirely unknown and unquantified (Rising et al., 2022). As quantitative and qualitative methodologies for assessing such risks continue to evolve, their integration into DFA frameworks may become more feasible.
@@ 2.2.1 Climate module
    Bias correction: For each ensemble member, we correct bias by comparing model backcasts to historical observations via the quantile mapping method, a simple but effective technique frequently used in the literature (Haug et al., 2011; Maraun, 2013; Sanabria et al., 2022). Specifically, we estimate
@@
    where $\theta_{q}$ and $\hat{\theta}_{q}^{(m)}$ are the $q^{\text{th}}$ quantiles of the historical observations and model $m$ backcasts, respectively, over the same reference period.
@@ 2.2.2 Hazards module
where $\bm{\Theta}^{(i)}_{t}$ is a set of weather covariates for hazard type $i$, and $X_{t}^{(i),m}$ is the normalised catastrophe loss adjusted for both inflation and wealth exposure  [^1] The per-event normalised loss is defined as: $X_{t}^{(i),m}=\tilde{X}_{t}^{(i),m}\cdot\frac{\text{CPI}_{s}}{\text{CPI}_{t}}\cdot\frac{\text{Real GDP}_{s}}{\text{Real GDP}_{t}}$, which adjusts the nominal loss ($\tilde{X}_{t}^{(i)}$) for both the price level and total wealth in reference year $s$. Although our normalization technique is relatively simple, it aligns with conventional approaches in the literature (see, e.g., Pielke and Landsea, 1998; Vranes and Pielke Jr, 2009). While more granular factors could be incorporated into normalization (see, e.g., Crompton and McAneney, 2008; Pielke, 2021), our focus on future projections rather than historical trends – and on national-level rather than granular losses – supports the adoption of a more parsimonious approach.. The Poisson distribution aligns with common practice in modelling the frequency of hazard events (see, e.g., Jagger et al., 2008, 2011). Similarly, Log-Normal distribution is frequently used for modelling catastrophe losses due to its heavy-tailed nature (Kaufmann et al., 2001; McNeil et al., 2015), and it is also selected from a class of heavy-tailed candidates based on our data (see Appendix A.3).
@@
$\log(\lambda^{(i)})=\bm{\beta}_{(i)}^{\prime}\bm{\Theta}_{t}^{(i)};\;\;\mu^{(i)}=\bm{\alpha}_{(i)}^{\prime}\bm{\Theta}_{t}^{(i)},$ (2.5)
where the set of coefficients $\bm{\beta}_{(i)}$ and $\bm{\alpha}_{(i)}$ are estimated via regression.
We select the weather covariates based on the physical mechanisms driving each hazard type $i$ and validate them statistically. Here, we focus on major Australian hazards: flood, bushfire, tropical cyclones, storms, hailstorm, and East Coast Lows (Insurance Council of Australia, 2024). Below, we outline the candidate weather covariates for each hazard type, drawn from relevant literature on the associated physical processes.
@@
where $\Delta\overline{\mathrm{SST}}_{m}$ represents the SST gradient near the east coast of Australia, defined as the average difference between $\overline{\mathrm{SST}}_{m}$ in the region spanning $24^{\circ}$$41^{\circ}\mathrm{S}$ and $148^{\circ}$$155^{\circ}\mathrm{E}$, and $\overline{\mathrm{SST}}_{m}$ in the corresponding region spanning $24^{\circ}$$41^{\circ}\mathrm{S}$ and $160^{\circ}$$165^{\circ}\mathrm{E}$ (Pepler et al., 2016a).
@@ 2.3.1 Inflation rates
$i_{m}^{\text{Clim-Impact}}=\sum_{L=1}^{11}(\alpha_{1+L}\Delta\bar{T}_{m-L}^{\text{NS}}+\beta_{1+L}\bar{T}_{m-L}^{\text{NS}}\cdot\Delta\bar{T}_{m-L}^{\text{NS}}),$ (2.8)
@@ 2.3.3 Equity module
$r_{t}^{(S)}=\tilde{r}(t)+x_{t},$ (2.11)
where $\tilde{r}(t)$ is the nominal risk-free rates, and $x_{t}$ is the excess equity return.
@@
To strike a balance between these two approaches, we propose a partial self-contained framework that incorporates forward-looking climate considerations without relying on overly granular external data. Inspired by the climate-economic literature (see, e.g., Karydas and Xepapadeas, 2022; Barnett, 2023), our method channels climate’s influence on equity returns through climate-damaged consumption. In our historical calibration, excess total equity returns are modelled as a function of corporate earnings growth, $\Delta\text{OP}_{t}$, which is itself modelled as a function of consumption growth, $\Delta C_{t}$. Specifically, we have:
$x_{t}=\beta_{0}+\beta_{1}\Delta\text{OP}_{t}+\epsilon_{t}^{\text{x}},\;\;\;\Delta\text{OP}_{t}=\alpha_{0}+\alpha_{1}\Delta C_{t}+\epsilon_{t}^{O},$ (2.12)
where $\epsilon_{t}^{\text{x}}\sim N(0,\sigma^{2}_{\text{x}})$ and $\epsilon_{t}^{O}\sim N(0,\sigma^{2}_{O})$, with the parameters calibrated via linear regression.
For future simulations, the following steps are undertaken. First, we obtain the simulated nominal market insurance catastrophe loss $\tilde{X_{t}}$ from the hazard module (see Section 2.2.2). Next, these insurance losses are scaled to represent uninsured economic damage, yielding $\eta\tilde{X}_{t}$. The production available for consumption after climate damage is then computed as:
$C_{t}=Y_{t}-\eta\tilde{X}_{t}.$ (2.13)
Subsequently, corporate earnings growth for the general sector is simulated based on Equation (2.12). For the brown sector, we apply a transition stress overlay factor (Grippa and Mann, 2020), which adjusts corporate earnings growth as:
@@
where $\Delta Y_{t}^{B}$ represents the change in brown energy production, and $\beta$ is the sensitivity of brown firms’ corporate profits to these changes. Finally, the total excess return level is derived based on the simulated operating profit growth using Equation (2.12).
Based on the outputs from the interest rate (Section 2.3.2) and equity modules, investment returns are calculated as: $r_{t}^{I}=w_{f}\tilde{r}_{t}+(1+w_{f})r_{t}^{(S)}$, where $w_{f}$ is the proportion of the portfolio allocated to risk-free assets.
In summary, the key inputs to this model are the nominal risk-free rate $\tilde{r}(t)$, aggregate catastrophe losses $\tilde{X}_{t}$, real GDP projections, and brown energy production for each SSP scenario ($Y_{t}$ and $Y_{t}^{B}$), with $\tilde{r}(t)$ and $\tilde{X}_{t}$ obtained from the respective interest rate and hazards modules. The model outputs are the equity returns for the general portfolio, $r_{t}^{(S,G)}$, and for the brown portfolio, $r_{t}^{(S,B)}$. By preserving the simplicity of traditional methods while integrating forward-looking climate damage projections, this partial self-contained approach captures key trends in evolving climate and socio-economic conditions without using high-dimensional external factors.

### v2

@@ 1.1 Background
Climate change poses a multifaceted and profound impact on the financial performance and position of general insurers. On the liabilities side, shifting weather patterns are likely to alter the frequency and severity of future claims (see, e.g., Haug et al., 2011; Lyubchich et al., 2019), which could lead to increased claims costs and adversely affect underwriting profitability. On the assets side, climate change-induced fluctuations in temperature, shifting patterns of natural disasters, and evolving socio-economic conditions may significantly influence key macroeconomic variables. These include inflation rates (see, e.g., Parker, 2018; Economides and Xepapadeas, 2018; Kotz et al., 2024), interest rates (see, e.g., Bylund and Jonsson, 2020; Mongelli et al., 2022), and equity returns (see, e.g., Karydas and Xepapadeas, 2022; Venturini, 2022; Barnett, 2023), thereby impacting the investment performance of general insurers. Ultimately, the combined effects of climate change on both assets and liabilities will cascade into insurers’ capital positions, potentially affecting the financial health and stability of the broader general insurance market.
@@
To navigate these challenges, a holistic understanding of climate change’s impact on general insurers is essential. This need is increasingly recognised by regulators, as evidenced by the growing number of climate-related disclosure requirements, such as IFRS S2 (IFRS, 2023) and its Australian equivalent, AASB S2 (AASB, 2024). In the existing literature, several studies have investigated the impacts of climate change on individual financial items. For example, on the liabilities side, various studies have examined climate change effects on hazards such as floods (see, e.g., Seneviratne et al., 2021; Bourget et al., 2024), bushfires (see, e.g., Quilcaille et al., 2022), and tropical cyclones and storms (see, e.g., Jagger et al., 2008, 2011; Meiler et al., 2022), generally indicating that rising temperatures exacerbate the frequency and severity of most hazard events (Chapter 11 of IPCC, 2021a, pp. 1546–1600). Additionally, the impact of climate change on non-catastrophe, weather-related claims frequency and severity have also been investigated in literature (see, e.g., Haug et al., 2011; Lyubchich et al., 2017). Outside the general insurance domain, the impacts of climate change on mortality rates have also been examined in the literature (Min et al., 2025; Guibert et al., 2025). On the assets side, climate change impacts on interest rates, inflation rates, and equity returns have been explored both theoretically using General Equilibrium Models (Economides and Xepapadeas, 2018; Karydas and Xepapadeas, 2022; Barnett, 2023) and empirically using multiple regression (see, e.g., Parker, 2018) and Factor Models (see, e.g., Venturini, 2022). While efforts have been made to incorporate climate change impacts into the modelling of insurance liabilities and assets in various special (and limited) cases, as yet a holistic approach to assess the interconnected financial impacts of climate change on general insurers across both assets and liabilities remains lacking.
@@ 1.2 Statement of contributions
    We also acknowledge that the impacts of climate change on individual risk types have been examined in various studies in the actuarial field. For example, Bourget et al. (2024) analyse the effect of climate change on pluvial flood occurrences. Outside the general insurance domain, Min et al. (2025) and Guibert et al. (2025) investigate climate change’s implications for mortality risk. Our paper extends those studies by providing a holistic modelling framework that assesses the financial impact of climate change on general insurers across their income statements, integrating multiple risk types and capturing the dependencies between them.
@@
    Long-term nature of climate change: Climate change is inherently long-term, and it is usually challenging to distinguish the effects of global warming from internal variability over short horizons; indeed, internal variability ranges from inter-annual (e.g., El Niño–Southern Oscillation) to inter-decadal (e.g., Atlantic Multi-decadal Variability) (Chapter 11 of IPCC, 2021a, pp. 1546–1600). In addition, new regulations – such as IFRS S2 (IFRS, 2023) and AASB S2 (AASB, 2024) – require insurers to disclose the long-term financial impacts of climate change. A longer-term perspective is also essential for decisions by insurers and policymakers that extend across multiple years, including relocation planning (typically 8 to 40 years Bower and Weerasinghe, 2021) and reinsurance planning (influenced by cycles of about 9 to 11 years Meier and Outreville, 2006, 2010).
    An initial attempt to measure climate change effects on both assets and liabilities of general insurers simultaneously is presented in Gatzert and Özdil (2024), albeit limited to a one-year horizon. Our proposed framework utilizes a multi-year perspective within a DFA structure to assess the cumulative effects of climate change on general insurers. Unlike most solvency frameworks, where the required assessment is restricted to a one-year time horizon (e.g., Chapter 2, Section 2 of Delegated Regulation (EU) 2015/35 (EIOPA, 2015), Chapter 8 of SS17/16 (PRA, 2016)), a multi-year framework is essential for evaluating climate change impacts, given their inherently long-term nature.
@@
    By enabling simulation of a distribution of financial outcomes across diverse climate pathways under the DFA setting, the framework supports assessments of both expected returns and risks, which are key considerations for insurers, actuaries, and regulators. While our analysis incorporates the economic and environmental assumptions underlying each scenario, evaluating the validity of those assumptions lies beyond the scope of this paper.
Our proposed framework is tailored to the general insurance sector by incorporating its unique features: it captures asset–liability interdependence through an interconnected structure and models the high variability of liability cash flows via DFA’s stochastic simulations. It also includes catastrophe reinsurance programs and accounts for the sensitivity of reinsurance premiums to capital constraints, which are also key factors affecting insurer profitability and solvency (see, e.g., Meier and Outreville, 2006). This differs from studies such as Min et al. (2025) and Guibert et al. (2025), which analyse climate change impacts on mortality risk, a risk type more pertinent to life insurance.
@@ 2 Modelling framework for climate-dependent DFA: A DFA framework underpinned by climate inputs
The framework is designed to be general, with most component models described in Sections 2.2 to 2.5 being transferable to other insurance markets. However, while the overall structure of the hazard module is adaptable, the specific hazard types and climate covariates must be tailored to the region under study. In Section 2.2, we illustrate the modelling of individual hazard types and the selection of relevant climate covariates by using Australia as an example, while offering generalised guidance where applicable.
@@ 2.1 Model overview
The proposed climate-dependent DFA framework is illustrated in Figure 2.1. A key input to this framework is the set of climate scenarios, for which we adopt the Shared Socioeconomic Pathways (SSPs). These SSPs form a widely adopted framework in climate research and they are central to the IPCC’s climate risk assessments (O Neill et al., 2017). It should be noted, however, that the SSPs represent only one of several possible scenario frameworks. In particular, it is acknowledged that the SSPs may be succeeded by the Representative Emission Pathways (REPs) in the IPCC’s Seventh Assessment Report (Meinshausen et al., 2024), which aim to provide more policy-relevant, flexible, and up-to-date scenario representations. The integration of such alternative scenarios is left for future research.
Each SSP scenario is associated with a narrative, from which the economic growth rate at the technological frontier is derived (Dellink et al., 2017a). Starting from historical values, country-specific GDP projections are generated under the assumption that individual economies gradually converge toward this frontier. The convergence speed is determined by the degree of trade openness, as inferred from the scenario narratives (Dellink et al., 2017a). The emissions pathways consistent with the economic and environmental assumptions underlying each scenario are then used as inputs to climate models to produce projections of future climate at a much finer spatial resolution, typically at the level of gridded cells (Eyring et al., 2016).
@@
    SSP 4.5 (“Middle of the Road”)  [^1] We follow the naming convention for the SSP 4.5 scenario as used in the literature (O Neill et al., 2017). However, it should be noted that projections under SSP 4.5 might underestimate both emissions and temperature increases under current policies. For instance, while the best estimate of global temperature increase under SSP 4.5 is about 1.5 degree by 2030, this level of warming had already been reached by 2023 (Goessling et al., 2025).: Represents a development pathway aligned roughly with typical historical trends observed over the past century, leading to moderate mitigation and adaptation challenges.
@@
Note that caution is warranted when interpreting the results derived from these scenarios, given the limitations of their underlying assumptions, particularly under high-emission pathways such as SSP 8.5. This scenario assumes continued economic growth without accounting for the risk of economic collapse (Keen et al., 2021). Moreover, labour productivity is treated as exogenous in SSP scenarios, overlooking potential negative impacts of climate change (Keen et al., 2021). In addition, recent studies suggest that the increasing duration and intensity of extreme heat under high-emission scenarios tend to reduce life expectancy (Guibert et al., 2025; Min et al., 2025), which could, in turn, negatively affect labour force growth. However, this potential impact is likewise not incorporated into the labour force growth assumptions in the current SSP frameworks. While SSP 8.5 also assumes strong investments in health, education, and highly engineered infrastructure (O Neill et al., 2017), which may mitigate climate-related productivity and labour force losses, the extent of such mitigation remains uncertain. In addition, the SSP scenarios do not capture abrupt or disorderly transition pathways and are therefore less suitable for stress tests involving severe transition shocks to the economy. Assessing or refining these assumptions lies beyond the scope of this paper and is left for future research.
@@
This paper focuses on the direct financial impacts of climate change on general insurers, as outlined in the modelling framework presented in Figure 2.1. While indirect effects, such as shifts in customer preference toward “greener” insurers or the potential rise in liability risks (e.g., lawsuits against commercial policyholders for environmental damage Alien, 2003; Bullock, 2022) that represent transition risks on the liability side are not included here, these are important areas for future research. In addition, we do not consider unidentified risks stemming from unforeseen responses of environmental and social systems to climate change, as these currently remain entirely unknown and unquantified (Rising et al., 2022). As quantitative and qualitative methodologies for assessing such risks continue to evolve, their integration into DFA frameworks may become more feasible.
@@ 2.2.1 Climate module
    Bias correction: For each ensemble member, we correct bias by comparing model backcasts to historical observations via the quantile mapping method, a simple but effective technique frequently used in the literature (Haug et al., 2011; Maraun, 2013; Sanabria et al., 2022). Specifically, using a quantile mapping approach with a linear transformation function (Piani et al., 2010; Qian and Chang, 2021), we estimate
@@
    where $\hat{\theta}_{q}$ and $\hat{\theta}_{q}^{(m)}$ are the $q^{\text{th}}$ quantiles of the historical observations and model $m$ backcasts, respectively, over the same reference period.
@@ 2.2.2 Hazards module
where $\bm{\Theta}^{(i)}_{t}$ is a set of weather covariates for hazard type $i$, and $X_{t}^{(i),m}$ is the normalised catastrophe loss adjusted for both inflation and wealth exposure  [^2] The per-event normalised loss is defined as: $X_{t}^{(i),m}=\tilde{X}_{t}^{(i),m}\cdot\frac{\text{CPI}_{s}}{\text{CPI}_{t}}\cdot\frac{\text{Real GDP}_{s}}{\text{Real GDP}_{t}}$, which adjusts the nominal loss ($\tilde{X}_{t}^{(i)}$) for both the price level and total wealth in reference year $s$. Although our normalization technique is relatively simple, it aligns with conventional approaches in the literature (see, e.g., Pielke and Landsea, 1998; Vranes and Pielke Jr, 2009). While more granular factors could be incorporated into normalization (see, e.g., Crompton and McAneney, 2008; Pielke, 2021), our focus on future projections rather than historical trends – and on national-level rather than granular losses – supports the adoption of a more parsimonious approach.. The Poisson distribution aligns with common practice in modelling the frequency of hazard events (see, e.g., Jagger et al., 2008, 2011). Similarly, Log-Normal distribution is frequently used for modelling catastrophe losses due to its heavy-tailed nature (Kaufmann et al., 2001; McNeil et al., 2015), and it is also selected from a class of heavy-tailed candidates based on our data. The detailed discussion of the selected distributions based on the Australian data is provided in Online Appendix A.3.1. However, alternative distributions may provide a better fit for other datasets; hence, it is recommended to validate the distributional assumptions proposed here when applying the model to new data.
@@
$\log(\lambda^{(i)}_{t})=\bm{\beta}_{(i)}^{\prime}\bm{\Theta}_{t}^{(i)};\;\;\mu^{(i)}_{t}=\bm{\alpha}_{(i)}^{\prime}\bm{\Theta}_{t}^{(i)},$ (2.5)
where the set of coefficients $\bm{\beta}_{(i)}$ and $\bm{\alpha}_{(i)}$ are estimated via regression. We do not explicitly adjust the scale parameter of the severity distribution in this study. However, it could be modelled as a function of climate covariates using GAMLSS regression (Stasinopoulos and Rigby, 2008). This is left for future research when stronger statistical and climate science evidence becomes available on the influence of climate covariates on the scale parameter.
We select the weather covariates based on the physical mechanisms driving each hazard type $i$ and validate them statistically. Here, we focus on major Australian hazards: flood, bushfire, tropical cyclones, storms, hailstorm, and East Coast Lows (Insurance Council of Australia, 2024). Below, we outline the candidate weather covariates for each hazard type, drawn from relevant literature on the associated physical processes. Although the data primarily cover Australia, the selection of candidate climate variables can be generalised to other regions, except for East Coast Lows, which is a unique hazard type occurring near the eastern coast of Australia.
The hazard model presented in this section is designed to capture industry-level trends consistent with the scope of this paper. For more granular decision-making at the level of individual insurers (e.g., portfolio management), a higher-resolution modelling framework may be required. A discussion of how the current hazard model can be extended to regional projections and integrated with CAT model outputs is provided in Online Appendix E. Nonetheless, incorporating such extensions should be carefully balanced against the trade-off between precision and interpretability depending on the intended business applications, as discussed in Section 1.2.
@@
where $\Delta\overline{\mathrm{SST}}_{m}$ represents the SST gradient near the east coast of Australia, defined as the average SST anomaly in the region $24^{\circ}$$41^{\circ}\mathrm{S}$ and $148^{\circ}$$155^{\circ}\mathrm{E}$ relative to the SST at the same latitudes and time between $160^{\circ}$$165^{\circ}\mathrm{E}$ (Pepler et al., 2016a).
@@ 2.3.1 Inflation rates
$i_{m}^{\text{Clim-Impact}}=\sum_{L=0}^{11}(\alpha_{1+L}\Delta\bar{T}_{m-L}^{\text{NS}}+\beta_{1+L}\bar{T}_{m-L}^{\text{NS}}\cdot\Delta\bar{T}_{m-L}^{\text{NS}}),$ (2.8)
@@ 2.3.3 Equity module
$r_{t}^{(S)}=\tilde{r}_{t}+x_{t},$ (2.11)
where $\tilde{r}_{t}$ is the nominal risk-free rates, and $x_{t}$ is the excess equity return.
@@
To strike a balance between these two approaches, we propose a partial self-contained framework that incorporates forward-looking climate considerations without relying on overly granular external data. Inspired by the climate-economic literature (see, e.g., Karydas and Xepapadeas, 2022; Barnett, 2023), our method channels climate’s influence on equity returns through climate-damaged consumption. A flow diagram illustrating the simulation of excess equity returns is presented in Figure 2.3.
Figure 2.3: Illustrative diagram showing the simulation flow for equity excess returns
In the first step, we obtain the simulated nominal market insurance catastrophe loss $\tilde{X_{t}}$ from the hazard module (see Section 2.2.2). Next, these insurance losses are scaled to represent uninsured economic damage, yielding $\eta\tilde{X}_{t}$, where $\eta$ is the multiplier. The production available for consumption after climate damage is then computed as:
$C_{t}=Y_{t}-\eta\tilde{X}_{t},$ (2.12)
where $Y_{t}$ is the real GDP projection under each SSP scenario.
To transmit the effects of economic growth and climate damage to equity returns, excess total equity returns are modelled as a function of corporate earnings (based on operating profits) growth, $\Delta\text{OP}_{t}$, which is itself modelled as a function of consumption growth, $\Delta C_{t}$. Specifically, we have:
$\Delta\text{OP}_{t}=\alpha_{0}+\alpha_{1}\Delta C_{t}+\epsilon_{t}^{O},\;\;\;x_{t}=\beta_{0}+\beta_{1}\Delta\text{OP}_{t}+\epsilon_{t}^{\text{x}}$ (2.13)
where the parameters are estimated through regression using historical data. The random shocks to operating profit growth and excess equity returns are represented by $\epsilon_{t}^{\text{x}}\sim N(0,\sigma^{2}_{\text{x}})$ and $\epsilon_{t}^{O}\sim N(0,\sigma^{2}_{O})$, with their volatilities calibrated from historical observations.
For the brown sector, we further apply a transition stress overlay factor (Grippa and Mann, 2020) on its operating profit growth:
@@
where $\Delta Y_{t}^{B}$ represents the change in brown energy production, and $\beta$ is the sensitivity of brown firms’ corporate profits to these changes. The adjusted operating profit growth for the brown portfolio is then used to simulate its excess equity returns based on (2.13).
Based on the outputs from the interest rate (Section 2.3.2) and equity modules, investment returns are calculated as: $r_{t}^{I}=w_{f}\tilde{r}_{t}+(1-w_{f})r_{t}^{(S)}$, where $w_{f}$ is the proportion of the portfolio allocated to risk-free assets.
In summary, the key inputs to this model are the nominal risk-free rate $\tilde{r}_{t}$, aggregate catastrophe losses $\tilde{X}_{t}$, real GDP projections, and brown energy production for each SSP scenario ($Y_{t}$ and $Y_{t}^{B}$), with $\tilde{r}_{t}$ and $\tilde{X}_{t}$ obtained from the respective interest rate and hazards modules. The model outputs are the equity returns for the general portfolio, $r_{t}^{(S,G)}$, and for the brown portfolio, $r_{t}^{(S,B)}$. By preserving the simplicity of traditional methods while integrating forward-looking climate damage projections, this partial self-contained approach captures key trends in evolving climate and socio-economic conditions without using high-dimensional external factors.
@@
In addition, when deriving net production after accounting for climate damage, we do not explicitly incorporate the reconstruction effects of natural disasters on GDP. The direction and magnitude of these effects remain largely inconclusive in the empirical literature. Albala-Bertrand (1993) found that, for most countries in their sample, short-term GDP growth increased following natural disasters, primarily due to the replacement of destroyed capital with more efficient ones, while the long-term growth rate remained unaffected. Skidmore and Toya (2002) reported that climatic disasters can also exert a positive influence on long-run economic growth in addition to short-term effects. In contrast, based on cyclone impacts, Hsiang and Jina (2014) rejected the hypothesis that disasters stimulate growth, showing that national income tends to decline and fails to recover in the short term. Similarly, Noy (2009) identified statistically significant short-term GDP losses, but with smaller magnitudes in developed economies due to their greater capacity to mobilize reconstruction resources. The theoretical framework of Hallegatte and Dumas (2009) further suggests that while disasters can affect the level of production, they do not alter the long-run growth rate, and the production response depends on the quality of reconstruction. Our assumption regarding economic damage shares some slight (and limited) similarity with Hallegatte and Dumas (2009) and Noy (2009), implying a short-run decline in production but no long-term effect on economic growth. However, the explicit incorporation and testing of alternative assumptions regarding the reconstruction effects on GDP are left for future research.